plot a graph of voltage against current and find the slope, the slope you find from voltage to current is the resistance which is measured in ohm's.


the standard error on slope is given as 4(V2-V1)/6(12-11)

1
Expert's answer
2022-01-17T10:05:49-0500

SAMPLE RESULTS


"\begin{matrix}
 Current (A)& &0.1&0.2&0.3&0.4 \\
 Voltage (V)& &1.2&2.4&3.6&4.8
\end{matrix}"








Slope of the graph


"=\frac {\Delta V}{\Delta I}"


"=\frac{(4.2-1.0)}{(0.35-0.085)}=12.075\Omega"


As per equation given

Standard error"=\frac{4(V_2-V_1)}{6(I_2-I_1)}"

"=\frac{4}{6}×slope"


"=\frac{4}{6}×12.075"

"=8.05"